Gucchi:

Compare the slopes of the linear functions f(x) and g(x) and choose the answer that best describes them. The slope of f(x) is greater than the slope of g(x). The slope of f(x) is less than the slope of g(x). The slope of f(x) is equal to the slope of g(x). The slope of g(x) is undefined.

Gucchi:

1 attachment
Gucchi:

i need help please

Gucchi:

@carmelle

Gucchi:

@axie

Gucchi:

@tetsxpreme

carmelle:

do you know the equation for slope?

Gucchi:

y = mx + b or the y2 - y1/ x2-x1 equation?

carmelle:

y2 - y1/ x2-x1

Gucchi:

yes ofc

Gucchi:

i remember

carmelle:

so what are you 2 points for ur graph?

Gucchi:

uh (-3, -2) and

carmelle:

what? where did that point come from? pick a point from the line

Anonumous300:

Its B

Gucchi:

wait the points are already on the graph right?

Gucchi:

@carmelle wrote:
what? where did that point come from? pick a point from the line
-1?

Anonumous300:

because f(x) is 1 up and 1 1/2 over and g(x) is 3 up and 2 over

Gucchi:

@anonumous300 wrote:
because f(x) is 1 up and 1 1/2 over and g(x) is 3 up and 2 over
thats true

carmelle:

Here are 2 points u can do

Gucchi:

ok so basically 2/3?

Gucchi:

because 4/6 simplifies

carmelle:

yep! how about the table?

Gucchi:

ok so lemme do the equation for the table

Gucchi:

is the y g(x)?

carmelle:

yes, g(x) is another way of saying y

Gucchi:

3/2 srry it took so long had to do sum

carmelle:

correct, so is 3/2 greater than, less than, or equal to 2/3

Gucchi:

Ok, so b

carmelle:

yep, that would be correct :)

carmelle:

\[\frac{ 2 }{ 3 } < \frac{ 3 }{ 2 }\]